OpenStudy (anonymous):

The polynomial of degree 3, P(x), has a root of multiplicity 2 at x=5 and a root of multiplicity 1 at x=-4. The y-intercept is y= -50 . Find a formula for P(x) .

OpenStudy (anonymous):

f(x) = (x-5)^2(x--4) = (x-5)^2(x+4) f(x) = x^2-5x-5x-25=(x^2-10x-25)(x+4) f(x) = x^3+4x^2-10x^2-40x-25x-100 = x^3-6x^2-65x-100 What do I do with the y intercept?

OpenStudy (anonymous):

I also wonder if I did too much work for this problem

OpenStudy (anonymous):

Just got it, thanks.